Three towns A, B, C are situated such that /AB/=25km and /AC/=30km. The bearing of B from A is 56°and bearing of C from A is 282°. Calculate the distance /BC/


1
Expert's answer
2021-11-24T11:59:09-0500

"360\u00b0-282\u00b0=78\u00b0"


"78\u00b0+56\u00b0=134\u00b0"


"BC=\sqrt{25^2+30^2-2\cdot25\cdot30\cdot\cos134°}=50.7\ (km)" . Answer